OpenStudy (anonymous):

find the length of the arc on a circle with a radius of 30 ft that is intercepted by a central angle measuring 2 1/2 radians

OpenStudy (anonymous):

2 1/2*30 ft=?

OpenStudy (anonymous):

is that all you do?

OpenStudy (anonymous):

no... what is the total circumference?

OpenStudy (anonymous):

360

OpenStudy (anonymous):

C = 2*pi*r

OpenStudy (anonymous):

so would it be... 360=2*pi*30? do I put the 360?

OpenStudy (anonymous):

you don't... there are 360 degrees in a circle. there are 2*pi radians in a circle too. these are equivalent. the arc of the circle encompassed by 2 1/2 radians is proportional to the total \[\frac{ 2\frac{ 1 }{ 2 } }{2\pi } \text{ is the proportion of the circle's circumference covered by this \angle}\] so \[\text{Arc length }=\left(\frac{2\frac{1}{2}}{2\pi}\right)2\pi r =2\frac{1}{2}\pi r \]

OpenStudy (anonymous):

thank you so much for your help! you made it a lot easier to understand :) could you please help me with 1 more problem?

OpenStudy (anonymous):

yeah... what is it?

OpenStudy (anonymous):

oops... i forgot to cancel the pi... it should be \[\text{Arc length }=2\frac{1}{2} r \]

OpenStudy (anonymous):

okay thank you! and its, find the central angle of a sector in degrees if the area of the sector is 220 ft^2 and the radius of the circle is 12 feet.

OpenStudy (anonymous):

So what's the area of a sector?

OpenStudy (anonymous):

is it A=a/2*r^2?

OpenStudy (anonymous):

yes, if a = angle in radians... so plug in what you know and solve for what you need...

OpenStudy (anonymous):

if you want, you can do that witthout numbers and then plug them in... \[ A = \frac{\theta}{2}r^{2} \text{ solve for }\theta\]

OpenStudy (anonymous):

so 220=a/2 (12)^2?

OpenStudy (anonymous):

yeah, now solve for a...

OpenStudy (anonymous):

do you divide by 24?

OpenStudy (anonymous):

I mean 144

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so is the answer 3.05 repeated?

OpenStudy (anonymous):

\[ 220=\frac{\theta}{2}(12)^{2}=\frac{\theta}{2}(144)=72\,\theta\Rightarrow \theta=\frac{220}{72}=\frac{55}{18}\]

OpenStudy (anonymous):

thank you for having the patience for helping me! :)

OpenStudy (anonymous):

thanks for persisting and learning!

OpenStudy (anonymous):

:)